# A relativistic model of b-EMRI systems and their gravitational radiation

### v1

@@ Abstract
A binary extreme-mass-ratio inspiral (b-EMRI) is a hierarchical triple system consisting of a stellar-mass binary black hole (BBH) orbiting a central Kerr supermassive black hole (SMBH). Although predicted by several astrophysical models, b-EMRIs pose a challenge in waveform modeling due to their complex three-body dynamics and strong relativistic effects. Here we take advantage of the hierarchical nature of b-EMRI systems to transform the internal motion of the small binary into global trajectories around the SMBH. This allows us to use black hole perturbation theory to calculate both the low-frequency gravitational waveform due to its EMRI nature and the high-frequency waveform generated by the inner motion of the BBH. When the inner binary’s separation vanishes, our calculation recovers the standard relativistic adiabatic EMRI waveform. Furthermore, by including the high-frequency perturbation, we find a correction to the waveform as large as the adiabatic order when the frequency matches the quasinormal modes (QNMs) of the SMBH, therefore supporting an earlier proof-of-concept study claiming that the small BBH can resonantly excite the QNMs of the SMBH. More importantly, we find that b-EMRIs can evolve faster than regular EMRIs due to this resonant dissipation through the high-frequency modes. These characteristics distinguish b-EMRI waveform templates from regular EMRI templates for future space-based gravitational-wave detectors.
@@ I Introduction
Extreme-mass-ratio inspirals (EMRIs) are an important target for future milli-Hertz gravitational-wave (GW) detectors, such as the Laser Interferometer Space Antenna (LISA) [1, 2], or China’s Taiji [3] and TianQin [4, 5] projects. An EMRI normally consists of a stellar-mass black hole (BH) or neutron star (the ‘secondary’) and a SMBH (the ‘primary’). The mass ratio $\epsilon$ of the two bodies falls in a typical range between $10^{-6}$ and $10^{-3}$. By studying the GWs radiated during the last $\mathcal{O}(1/\epsilon)$ orbits before plunge [6], we can measure the spin and mass of the central SMBH, and test the Kerr metric with unprecedented precision [7, 8, 9, 10, 11]. There has been significant efforts in developing both approximate and fully relativistic EMRI waveform models over the years, e.g. [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], including many works on the finite size effects of the compact secondary body [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Recent astrophysical models predict that the small body in an EMRI, in fact, could be a stellar-mass BBH. Such a BBH could be either captured by the SMBH during a close tidal interaction [39, 40], or delivered to the vicinity of the SMBH by an accretion disk if the SMBH happens to be inside an active galactic nucleus (AGN) [41, 42, 43, 44]. In such a triple system, the BBH is often referred to as the “inner binary” (IB), and the binary formed by the IB and the SMBH is called the “outer binary” (OB). Since the small BH binary is now spiraling into the SMBH, the whole system is now known as the “binary EMRI” (or b-EMRI for short [40, 45]).
@@
While solid observational evidence for b-EMRIs is currently lacking, it has been suggested that the event GW190521, discovered in the third observing run of the Laser Interferometer GW Observatory (LIGO) and the Virgo detectors [46, 47], could have derived from a BBH merger in vicinity of a SMBH. In particular, this event involves one of the most massive BBHs ever detected, whose final mass is approximately $150M_{\odot}$, far greater than the typical mass of the BHs found in Galactic X-ray binaries [48, 49]. The high mass could be explained by a redshift of the GW signal if the merger happened very close to a SMBH [50]. Moreover, a potentially associated AGN flare is reported by the Zwicky Transient Facility [51] and further analysis of the flare [52] suggests that GW190521 could come from the accretion disk of that AGN.
@@
The possibility of forming b-EMRIs in astrophysical environments has inspired many studies on the waveform of such a special GW source. The methods adopted by these earlier studies can be broadly divided into two categories. The first approach starts from the waveform of an isolated stellar-mass BBH and adds to it relativistic effects to account for the existence of a nearby SMBH [53]. Previously considered effects include Doppler and gravitational redshifts [54, 55, 56, 57, 58], aberration due to the fast motion of the BBH [59, 60], and gravitational lensing of the GW signal by the SMBH either in the geometric optics limit [61] or the wave optics one [62]. An important ingredient that is missing by these works is the gravitational radiation of the OB, i.e., the EMRI signal.
@@
Typical EMRI models employ a multi-scale (or two-timescale) expansion [6, 82, 83]. Within the multi-scale framework, the resulting waveforms are labelled by the accuracy of the waveform phase. This follows from the foundational scaling argument in Ref. [6], that showed that the average waveform phase has the asymptotic small mass-ratio expansion
@@ II.1 Definition of model parameters
Fig. 1 shows the physical picture of a b-EMRI system. The central SMBH is a Kerr black hole with a mass of $M\sim 10^{5}-10^{8}M_{\odot}$ and a spin parameter of $a$. We denote $p$ as the semi-latus rectum of the OB, which is equivalent to the radius of the OB if the orbit is circular. As for the IB, we denote the masses of the two stellar-mass BHs as $m_{1}$ and $m_{2}$. We use $\mu$ to represent the total mass of the IB so that $\mu\equiv m_{1}+m_{2}=M\epsilon$, and $m$ is reserved to denote a multipole number, which is an integer. We assume $m_{1}=m_{2}$ and a circular orbit for the IB for simplicity and denote its diameter as $d$.
Figure 1: A b-EMRI system and its parameters.
In the following, we will assume a Newtonian orbit for the inner orbital motion of the two stellar BHs, which requires that $d$ is relatively large compared to $\mu$ (e.g., $d>1000\mu$). Other constraints on $d$, $\epsilon$, and $p$ will be discussed in detail in Sec. V.2. There are two Euler angles determining the orientation of the IB and we will see how they function in Eq. (7), where $\tilde{\psi}$ is the initial precession angle, which introduces a phase term in the waveform, and $\tilde{\theta}$ is the inclination angle between the inner orbital plane and the outer one. We will see that $\tilde{\theta}$ is relatively more important than $\tilde{\psi}$ in affecting the b-EMRI waveform. In total there are seven parameters in our b-EMRI model. They are listed in Table 1.
@@
$\epsilon$ mass ratio $\sim 10^{-6}-10^{-3}$
@@
$\tilde{\psi}$ precession angle ranging from 00 to $2\pi$
$\tilde{\theta}$ inclination angle ranging from $-\pi/2$ to $\pi/2$
Table 1: Key parameters of b-EMRI in this paper.
@@ II.2 Orbital motion of the outer binary
In this study we assume that the orbit of the OB follows a circular-equatorial orbit in Kerr spacetime, the metric of which is
@@
where $\Delta=r^{2}-2Mr+a^{2}$ and $\Sigma=1/(\rho\bar{\rho})$ with $\rho=(r-ia\cos\theta)^{-1}$. Here overbar denotes complex conjugate. The complete form of the Kerr geodesics can be found in Ref. [84, 85, 86], but here we focus on the circular-equatorial case, where the Carter constant $\mathcal{Q}$ is set to zero. We express the geodesic equations in terms of the four-velocity $u^{x}=\mathrm{d}x/\mathrm{d}\tau\ (x=t,r,\theta,\varphi)$ as follows
@@ II.3 Coordinate transformation for the inner orbit
The relative motion of the two stellar-mass BHs can be solved most easily in a frame in which the binaries centre of mass is freely falling in the curved spacetime around the BBH (the aforementioned FFF). In this frame, the spacetime is sufficiently flat due to Einstein’s Equivalence Principle, and hence the two BH follows a near-Keplerian orbit. Together with the previous assumption of a circular IB, we can calculate the inner orbital frequency as $\omega_{IB}=\sqrt{\mu/d^{3}}$ and write the coordinates of the two BHs in their orbital plane as
$\tilde{x}=\frac{1}{2}d\cos(\omega_{IB}\tau),$ (6a)
@@
$\tilde{z}=\frac{1}{2}d\sin(\omega_{IB}\tau),$ (6c)
@@
$\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}\cos\tilde{\psi}&0&-\sin\tilde{\psi}\\ 0&1&0\\ \sin\tilde{\psi}&0&\cos\tilde{\psi}\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&\cos\tilde{\theta}&-\sin\tilde{\theta}\\ 0&\sin\tilde{\theta}&\cos\tilde{\theta}\end{bmatrix}\begin{bmatrix}\tilde{x}\\ \tilde{y}\\ \tilde{z}\end{bmatrix},$ (7)
to rotate the orbital plane of the IB and get the coordinates in the FFF.
Since the Teukolsky formalism is derived in BL coordinates, we need to transform the motion of the small BHs from the FFF to BL coordinates. This transformation is done in two steps [87, 80]. The relationship between different coordinates are illustrated in Fig. 2.
Figure 2: Different reference frames used in this paper and their relative orientation as they rotate around the SMBH. It is a top view of Fig. 1 where $e_{1},e_{2},e_{3}$ are unit bases with direction aligned with $e_{X^{1}},e_{X^{2}},e_{X^{3}}$.
First, we transform the inner motion into the LIF. This frame is moving at the same velocity as the FFF but its three spatial axes are aligned, respectively, with the $\hat{e}_{r}$, $\hat{e}_{\theta}$, and $\hat{e}_{\varphi}$ axes of the BL coordinates. It has been shown that the spatial axes of the FFF and LIF differ by a precession, with an angular frequency of $\Omega=-\sqrt{M/p^{3}}$ [87]. It happens to be the Keplerian orbital frequency in our circular-equatorial case. Consequently, coordinates in these two frames also differ by a rotation
@@
Second, we transform the coordinates in the LIF into the BL coordinates. Although the distance between each stellar-mass BH and the CM of the IB is fixed in the LIF (due to our assumption of circular inner orbit), in the BL coordinates the distance varies with time because of two reasons. (i) The coordinates in the LIF are measured by an observer at the CM of the IB, while the BL coordinates are measured by an observer at infinity. There is a coordinate transformation between the two. (ii) Since the CM is moving, the coordinate transformation also varies with time. It is important to mention that such a variation of the distance from the CM has not been accounted for in the previous studies. This varying distance can be generally expanded into an infinite series of terms [81], and for our purposes it is sufficient to truncate the terms at the second order in $X^{1,2,3}$. The resulting trajectory of one small BH has been derived in Ref. [80] and can be written as
@@
where the zeroth order terms $c_{00},\ c_{10},\ c_{20},$ and $c_{30}$ correspond to the outer motion with the geodetic relations:
@@
The first and second order coefficients $c_{01}$, $c_{02}$, $c_{11}$, $c_{12}$, $c_{21}$, $c_{22}$, $c_{31}$, $c_{32}$ are derived in Appendix A. We note that $c_{01}$, $c_{02}$, $c_{11}$, $c_{12}$, $c_{21}$, $c_{22}$, $c_{31}$, $c_{32}$ contain $\omega_{IB}=\sqrt{\mu/d^{3}}$, which is a function of $d$. However, we precalculate $\omega_{IB}$ here and treat it as another independent parameter.

### v2

# Relativistic model of binary extreme-mass-ratio inspiral systems and their gravitational radiation

@@ Abstract
A binary extreme-mass-ratio inspiral (b-EMRI) is a hierarchical triple system consisting of a stellar-mass binary black hole (BBH) orbiting a central Kerr supermassive black hole (SMBH). Although predicted by several astrophysical models, b-EMRIs pose a challenge in waveform modeling due to their complex three-body dynamics and strong relativistic effects. Here we take advantage of the hierarchical nature of b-EMRI systems to transform the internal motion of the small binary into global trajectories around the SMBH. This allows us to use black hole perturbation theory to calculate both the low-frequency gravitational waveform due to its EMRI nature and the high-frequency waveform generated by the inner motion of the BBH. When the inner binary’s separation vanishes, our calculation recovers the standard relativistic adiabatic EMRI waveform. Furthermore, by including the high-frequency perturbation, we find a correction to the waveform as large as the adiabatic order when the frequency matches the quasinormal modes (QNMs) of the SMBH, therefore supporting an earlier proof-of-concept study claiming that the small BBH can resonantly excite the QNMs of the SMBH. More importantly, we find that b-EMRIs can evolve faster than regular EMRIs due to this resonant dissipation through the high-frequency modes. These characteristics distinguish b-EMRI waveform templates from regular EMRI templates for future space-based gravitational wave detectors.
@@ I Introduction
Extreme-mass-ratio inspirals (EMRIs) are an important target for future milli-hertz gravitational wave (GW) detectors, such as the Laser Interferometer Space Antenna (LISA) [1, 2], or China’s Taiji [3] and TianQin [4, 5] projects. An EMRI normally consists of a stellar-mass black hole (BH) or neutron star (the ‘secondary’) and a SMBH (the ‘primary’). The mass ratio $\epsilon$ of the two bodies falls in a typical range between $10^{-6}$ and $10^{-4}$. By studying the GWs radiated during the last $\mathcal{O}(1/\epsilon)$ orbits before plunge [6], we can measure the spin and mass of the central SMBH, and test the Kerr metric with unprecedented precision [7, 8, 9, 10, 11]. There have been significant efforts in developing both approximate and fully relativistic EMRI waveform models over the years, e.g., [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], including many works on the finite-size effects of the compact secondary body [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Recent astrophysical models predict that the small body in an EMRI, in fact, could be a stellar-mass BBH. This type of BBH could be captured by the SMBH during a close tidal interaction [40, 41], or delivered to the vicinity of the SMBH by an accretion disk if the SMBH happens to be inside an active galactic nucleus (AGN) [42, 43, 44, 45]. In such a triple system, the BBH is often referred to as the “inner binary” (IB), and the binary formed by the IB and the SMBH is called the “outer binary” (OB). Since the small BBH is now spiraling into the SMBH, the whole system is now known as the “binary EMRI” (or b-EMRI for short [41, 46]).
@@
Although solid observational evidence for b-EMRIs is currently lacking, it has been suggested that the event GW190521, discovered in the third observing run of the Laser Interferometer GW Observatory (LIGO) and the Virgo detectors [47, 48], could have derived from a BBH merger in the vicinity of a SMBH. In particular, this event involves one of the most massive BBHs ever detected, whose final mass is approximately $150M_{\odot}$, much greater than the typical mass of the BHs found in Galactic X-ray binaries [49, 50]. The high mass could be explained by a redshift of the GW signal if the merger occurred very close to a SMBH [51]. Moreover, a potentially associated AGN flare is reported by the Zwicky Transient Facility [52] and further analysis of the flare [53] suggests that GW190521 could come from the accretion disk of that AGN.
@@
The possibility of forming b-EMRIs in astrophysical environments has inspired many studies on the waveform of such a special GW source. The methods adopted by these earlier studies can be broadly divided into two categories. The first approach starts from the waveform of an isolated stellar-mass BBH and adds to it relativistic effects to account for the existence of a nearby SMBH [54]. Previously considered effects include Doppler and gravitational redshifts [55, 56, 57, 58, 59, 60], aberration due to the fast motion of the BBH [61, 62], and gravitational lensing of the GW signal by the SMBH either in the geometric optics limit [63, 59] or the wave optics one [64]. An important ingredient that is missing by these works is the gravitational radiation of the OB, i.e., the EMRI signal.
@@
Typical EMRI models employ a multi-scale (or two-timescale) expansion [6, 84, 85, 39]. Within the multi-scale framework, the resulting waveforms are labelled by the accuracy of the waveform phase. This follows from the foundational scaling argument in Ref. [6], that showed that the average waveform phase has the asymptotic small mass-ratio expansion
@@ II.1 Definition of model parameters
Fig. 1a shows the physical picture of a b-EMRI system. The central SMBH is a Kerr black hole with a mass of $M\sim 10^{5}-10^{8}M_{\odot}$ and a spin parameter of $a$. We denote $p$ as the semi-latus rectum of the OB, which is equivalent to the radius of the OB if the orbit is circular. As for the IB, we denote the masses of the two stellar-mass BHs as $m_{1}$ and $m_{2}$. We use $\mu$ to represent the total mass of the IB so that $\mu\equiv m_{1}+m_{2}=M\epsilon$, and $m$ is reserved to denote a multipole number, which is an integer. We assume $m_{1}=m_{2}$ and a circular orbit for the IB for simplicity and denote its diameter as $d$.
Figure 1: A b-EMRI system and its parameters. Panel (a) illustrates the general scenario of a b-EMRI system. Panel (b) shows the Euler rotation between the inner orbital plane (in blue) and the outer orbital plane (in black). The blue arrows indicate the direction of motion of the small black holes. Note that the $y$-axis in panel (b) is geometrically aligned with the $e_{\theta}$-axis in the BL coordinates, which is opposite to the SMBH’s spin direction for equatorial OB.
In the following, we will assume a Newtonian orbit for the inner orbital motion of the two stellar BHs, which requires that $d$ is relatively large compared to $\mu$. However, if $d$ is too large, the IB will be unstable due to the Hill’s mechanism. More delicate constraints on $d$ will be provides in Sec. V.2. An appropriate choice is $d\approx\mbox{few thousand}\ \mu$ (see Sec. IV.1). For such a large $d$, we need a fairly small $\epsilon$ (e.g., $\epsilon<10^{-4}$) to maintain the stability of the system. Other constraints on $d$, $\epsilon$, and $p$ will also be discussed in detail in Sec. V.2.
As shown in Fig. 1b, there are three Euler angles determining the orientation of the IB and we will see how they function in Eq. (7). The $y$-axis of the FFF in Fig. 1b is naturally defined in the same direction as the $e_{\theta}$-axis of the BL coordinates in Fig. 1a, thus appearing downward when viewed in Fig. 1a. To facilitate our discussion of Euler rotation, we have reversed the system in Fig. 1b so that the spin direction of the SMBH is downward. However, the directions of the $x$- and $z$-axes do not coincide with those of $e_{r}$ and $e_{\varphi}$, but differ by a precession $\Omega$. To describe this precession, we define the LIF. The coordinate transformations between FFF, LIF, and BL will be thoroughly discussed in Sec. II.3. Form the perspective of Euler rotation, $\tilde{\psi}$ is the precession angle, $\tilde{\varphi}$ is the rotation angle and $\tilde{\theta}$ is the nutation angle. Physically, $\tilde{\theta}$ is also the angle between the inner orbital angular momentum and the outer one. The $\tilde{\varphi}$ can be absorbed into the inner motion, resulting in a degeneracy with the initial phase of the inner circular motion. We therefore set the initial phase of IB to zero to simplify the calculations. Besides, $\tilde{\psi}$ degenerates with the azimuthal angle of the observer, $\varphi$, in the BL coordinates. We therefore set $\tilde{\psi}$ to zero for the same reason. In total there are eight intrinsic parameters in our b-EMRI model. They are listed in Table 1.
@@
$\epsilon$ mass ratio $\sim 10^{-6}-10^{-4}$
@@
$\tilde{\psi}$ precession angle set to 00
$\tilde{\theta}$ inclination angle ranging from 00 to $\pi$
$\tilde{\varphi}$ rotation angle ranging from 00 to $2\pi$
Table 1: Key parameters of b-EMRI in this paper, where $M$, $a$, $\epsilon$, $p$, $d$ are shown in Fig. 1a, and $\tilde{\psi}$, $\tilde{\theta}$, $\tilde{\varphi}$ are shown in Fig. 1b
@@ II.2 Orbital motion of the outer binary
In this study, we assume that the orbit of the OB follows a circular-equatorial orbit in Kerr spacetime, the metric of which is
@@
where $\Delta=r^{2}-2Mr+a^{2}$ and $\Sigma=1/(\rho\bar{\rho})$ with $\rho=(r-ia\cos\theta)^{-1}$. Here overbar denotes complex conjugation. The complete form of the Kerr geodesics can be found in Ref. [86, 87, 88], but here we focus on the circular-equatorial case, where the Carter constant $\mathcal{Q}$ is set to zero. We express the geodesic equations in terms of the four-velocity $u^{x}=\mathrm{d}x/\mathrm{d}\tau\ (x=t,r,\theta,\varphi)$ as follows
@@ II.3 Coordinate transformation for the inner orbit
The relative motion of the two stellar-mass BHs can be solved most easily in a frame in which the binary’s CM is freely falling in the curved spacetime around the BBH (the aforementioned FFF). In this frame, the spacetime is sufficiently flat due to Einstein’s Equivalence Principle, and hence the two BH follows a near-Keplerian orbit. Together with the previous assumption of a circular IB, we can calculate the inner orbital frequency as $\omega_{\mathrm{IB}}=\sqrt{\mu/d^{3}}$ and write the coordinates of the two BHs in their orbital plane as
$\tilde{x}=\frac{1}{2}d\cos(\omega_{\mathrm{IB}}\tau),$ (6a)
@@
$\tilde{z}=\frac{1}{2}d\sin(\omega_{\mathrm{IB}}\tau),$ (6c)
@@
$\begin{bmatrix}x\\ y\\ z\end{bmatrix}=$ $\begin{bmatrix}\cos\tilde{\varphi}&0&\sin\tilde{\varphi}\\ 0&1&0\\ -\sin\tilde{\varphi}&0&\cos\tilde{\varphi}\end{bmatrix}\begin{bmatrix}\cos\tilde{\theta}&-\sin\tilde{\theta}&0\\ \sin\tilde{\theta}&\cos\tilde{\theta}&0\\ 0&0&1\end{bmatrix}$ (7)
×$\times\begin{bmatrix}\cos\tilde{\psi}&0&\sin\tilde{\psi}\\ 0&1&0\\ -\sin\tilde{\psi}&0&\cos\tilde{\psi}\end{bmatrix}\begin{bmatrix}\tilde{x}\\ \tilde{y}\\ \tilde{z}\end{bmatrix},$
to rotate the orbital plane of the IB and get the coordinates in the FFF. Note that the rotation of $\tilde{\varphi}$ can be absorbed as an initial phase in Eqs. (6), which has been set to zero in Eqs. (6).
Since the Teukolsky formalism is derived in BL coordinates, we need to transform the motion of the small BHs from the FFF to BL coordinates. This transformation is done in two steps [89, 82]. The relationships between different coordinates are illustrated in Fig. 2.
Figure 2: Different reference frames used in this paper and their relative orientation as they rotate around the SMBH. It is a top view of Fig. 1a where $e_{1},e_{2},e_{3}$ are unit vectors in the LIF, aligned with $e_{r},e_{\theta},e_{\varphi}$ in the BL coordinates. The FFF and the LIF share the same origin, centered on the CM of the BBH, while the origin of the BL coordinates is centered on the Kerr SMBH which is not shown in the figure.
First, we transform the inner motion into the LIF. This frame is moving at the same velocity as the FFF but its three spatial axes are aligned, respectively, with the $e_{r}$, $e_{\theta}$, and $e_{\varphi}$ axes of the BL coordinates. It has been shown that the spatial axes of the FFF and LIF differ by a precession, with an angular frequency of $\Omega=-\sqrt{M/p^{3}}$ [89]. It happens to be the Keplerian orbital frequency in our circular-equatorial case. There is a minus sign because the $e_{y}$ axis points downward when aligned with the $e_{\theta}$ axis in the equatorial plane. Consequently, coordinates in these two frames also differ by a rotational transformation
@@
Therefore, the $e_{x}$ and $e_{z}$ axes in FFF will coincide with the $e_{1}$ and $e_{3}$ axes in LIF periodically, i.e., at the moments (i) in Fig. 2. By comparing Fig. 2 (i) with Fig. 1a, we find that $\tilde{\psi}=0$ at these moments. We can choose these moments as reference points so that the initial precession information is contained in the azimuthal angle $\varphi$ between the BBH, the SMBH, and the observer. Therefore, the rotation with respect to $\tilde{\psi}$ in Eq. (7) becomes identical transformation.
Second, we transform the coordinates in the LIF into the BL coordinates. Although the distance between each stellar-mass BH and the CM of the IB is fixed in the LIF (due to our assumption of circular inner orbit), in the BL coordinates the distance varies with time because of two reasons. (i) The coordinates in the LIF are measured by an observer at the CM of the IB, while the BL coordinates are measured by an observer at infinity. There is a coordinate transformation between the two. (ii) Since the CM is moving, the coordinate transformation also varies with time. It is important to mention that such a variation of the distance from the CM has not been accounted for in the previous studies. This varying distance can be generally expanded into an infinite series of terms [83], and it is sufficient to truncate the terms at the second order in $X^{1,2,3}$ for our purposes. The resulting trajectory of one small BH has been derived in Ref. [82] and can be written as
@@
where the zeroth order terms $c_{00},\ c_{10},\ c_{20},$ and $c_{30}$ correspond to the outer motion with the geodesic relations:
@@
The first and second order coefficients $c_{01}$, $c_{02}$, $c_{11}$, $c_{12}$, $c_{21}$, $c_{22}$, $c_{31}$, $c_{32}$ are derived in Appendix A. We note that $c_{01}$, $c_{02}$, $c_{11}$, $c_{12}$, $c_{21}$, $c_{22}$, $c_{31}$, $c_{32}$ contain $\omega_{\mathrm{IB}}=\sqrt{\mu/d^{3}}$, which is a function of $d$. It is important to stress that we perform all of the series expansions in $d$ at fixed values of $\omega_{\mathrm{IB}}$ for several reasons. First of all, that is the most straightforward way to ensure the expansions are well-behaved when treating the disparate time and length scales of the outer and inner binaries. Secondly, it avoids inaccurate expansions of the oscillatory functions of the inner binary’s orbital phase. Finally, by holding $\omega_{\mathrm{IB}}$ fixed, we avoid cumbersome expansions of $d$-dependence of the homogeneous radial Teukolsky solutions via $\omega_{\mathrm{IB}}$ in the discrete frequency spectrum, as described in the following subsection. Thus, we precalculate $\omega_{\mathrm{IB}}$ and treat it as an independent parameter throughout. Its evolution is much slower than the outer binary’s evolution and, therefore, may be neglected.